OpenStudy (anonymous):

If the cost price of 20 articles is equal to the selling price of 25 articles, what is the % profit or loss made by the merchant?

OpenStudy (anonymous):

Let the SP of 1 article = x Then SP of 25 articles = 25x ATQ, CP of 20 articles = 25x CP of 1 article = 25x/20 = 5x/4 comparing x and 5x/4 we find 5x/4> x => CP is more than SP => loss

OpenStudy (anonymous):

Let the price be x.... so for 20 it becomes 20x now ur sayin that sp of 25 is sam to tht of 20 cp let sp be y 25y=20x it shows y = 4/5x or u can say 80% of x

OpenStudy (anonymous):

shopkeeper suffers a loss of 20% . so bad for the shopkeeper !

OpenStudy (anonymous):

sry merchant

OpenStudy (anonymous):

how 20%

OpenStudy (anonymous):

because y is 80% of x so .....(100 - 80)=20

OpenStudy (anonymous):

Loss Loss% = ------- x 100% CP 5x 5x - 4x x Loss = CP - SP = ----- - x = ---------- = ----- 4 4 4 x/4 x 4 1 Loss% = ------ x 00% = ----- * ------ * 100% = ----- * 100% = 20% 5x/4 4 5x 5

OpenStudy (anonymous):

Rupal, Is it clear ?????

OpenStudy (anonymous):

Loss of 20%....